# Optimizing wheel loader performance: an end-to-end approach

### v1

@@ Abstract
Wheel loaders in mines and construction sites repeatedly load soil from a pile to load receivers. This task presents a challenging optimization problem since each loading’s performance depends on the pile state, which depends on previous loadings. We investigate an end-to-end optimization approach considering future loading outcomes and V-cycle transportation costs. To predict the evolution of the pile state and the loading performance, we use world models that leverage deep neural networks trained on numerous simulated loading cycles. A look-ahead tree search optimizes the sequence of loading actions by evaluating the performance of thousands of action candidates, which expand into subsequent action candidates under the predicted pile states recursively. Test results demonstrate that, over a horizon of 15 sequential loadings, the look-ahead tree search is 6 % more efficient than a greedy strategy, which always selects the action that maximizes the current single loading performance, and 14 % more efficient than using a fixed loading controller optimized for the nominal case.
@@
[UMUAddress]organization=Umeå University,addressline=, city=Umeå, postcode=SE-90187, state=, country=Sweden
@@ 1 Introduction
The task of most construction and mining machines, such as wheel loaders, is to perform long sequences of earthmoving operations in dynamic environments. Optimizing this task involves selecting loading and transportation actions that maximize the gain over time, according to some performance measure. This is challenging for several reasons. First, the performance of a single loading cycle depends highly on the local shape of the pile and the soil properties [23]. The loading actions need to be carefully adapted to these conditions. Second, each loading alters the pile shape. From the perspective of future loadings, such a change can be either for the better or worse. There might be actions that are of low performance in the short term but are necessary to achieve higher gains in the future [22]. Therefore, a greedy strategy of always choosing the loading action that maximizes the performance for a single loading might be sub-optimal in the long run. Such a strategy might gradually deteriorate the pile state or lead to long transportation between the dig location and the load receiver. This motivates considering future states when selecting the actions.
@@
This paper investigates methods for solving the wheel loader end-to-end optimization problem and analyzes the obtained solutions. The research questions include what characterizes the optimal action sequence, how sensitive the results are to the initial pile state and planning horizon, what the computational demands are, and the feasibility of practical use of the method.
We use a look-ahead tree search approach illustrated in Fig. 1. Future action candidates for the foreseen future states are identified recursively. The action candidates are evaluated, using the data-driven models developed in previous research [2], to compute their respective performance and expand into further action candidates under the new states. This approach is, reportedly, similar to how experienced operators intuitively plan their work, but we expect a computerized model and search algorithms can make more accurate predictions and plan over longer horizons. An overview of the end-to-end optimization using a look-ahead tree search is illustrated in Fig. 1. To understand the importance of the look-ahead search, we examined the effect of using different planning horizons and different objectives.
@@
The wheel loader is assumed to have a high-level planner and a low-level control system, which work as the hierarchical architecture described in refs. [11] and [25]. The high-level planner selects a subtask action, such as bucket-filling, dumping, or V-cycle transportation, according to the current state of the machine and the work site. A low-level controller performs the subtasks given a motion plan and control parameters. For bucket filling, the machine is equipped with a type of admittance controller [6]. Our controller takes four action parameters that determine how reactive the boom and bucket actuation is in response to the momentaneous dig force [2]. The performance of a loading cycle is measured in terms of mass, time, and work, with contributions from each subtask.
@@ 2 Related work
What is a good dig location in the pile, given its shape in terms of a heightmap, has been studied from the perspective of both bucket filling and the transportation between the dig and the dump location. In [21], a coarse-level planner finds the points along the pile contour closest to the load receiver, and a refined planner is used to avoid curved regions of the pile surface to minimize bucket-side loading and drop-in filling efficiency. A sequential coarse-to-fine planner was studied in [14] and compared to a greedy planner. The coarse planner first produces a sequence of dig regions ordered radially around the pile centroid. The fine planner then selects the dig points inside each region that have the best local pile shape. Simulations (cellular automata) of 50 consecutive loadings showed that the coarse-to-fine strategy maintains a good pile shape and performs at 80–90 % while a greedy strategy (always selecting the dig location with optimal pile shape) initially performs equal or better for ten loadings but drops down to about 60 % after 25 loadings.
@@
The optimal bucket filling trajectory was researched in [7] using the discrete element method (DEM) and used to find the optimal short loading cycle using dynamic programming in [8]. For the sake of computational efficiency, the simulations were limited to quasi-2D piles sloped with the material’s angle of repose. The effect of the pile shape on the optimal trajectory or the resulting 3D pile state has not been explored using DEM simulation, apart from the work in [12].
@@ 3 Problem formulation
We consider the short loading cycle, where the wheel loader’s task is to load repeatedly soil from a pile onto a load receiver. Each cycle can be divided into a sequence of subtasks, V-turn-1, loading, V-turn-2, and dumping. They are illustrated in Fig. 1. We assume the wheel loader is equipped with a low-level control system for each subtask. A high-level system (operator or agent) activates the low-level systems by selecting a set of subtask action parameters: $\bm{a}^{\text{V1}}$, $\bm{a}^{\text{load}}$, $\bm{a}^{\text{V2}}$, and $\bm{a}^{\text{dump}}$, respectively.
At the beginning of each cycle $n$, the wheel loader is located at the load receiver at a position $\bm{x}_{n}^{\text{dump}}$. The V-turn-1 subtask is to drive the wheel loader to a selected loading position $\bm{x}_{n}^{\text{dig}}$ somewhere along the edge of the nearby pile. The low-level controller plans and executes a V-turn motion that starts at $\bm{x}_{n}^{\text{dump}}$ and ends at $\bm{x}_{n}^{\text{dig}}$, avoiding collisions with the environment and with endpoints heading normal to the load receiver and the pile, respectively. As subtask action parameters, we use the target loading position, $\bm{a}_{n}^{\text{V1}}=\bm{x}_{n}^{\text{dig}}$.
The next subtask is the actual loading from the pile. It starts with the machine approaching the target loading position and ends with the vehicle reversing from the pile with a filled bucket. A set of subtask action parameters, $\bm{a}_{n}^{\text{load}}$, is input to an automatic bucket filling controller. The loading transforms the pile from a state $\bm{H}_{n}$ into a new state $\bm{H}_{n+1}$ and results in some amount of soil mass $M_{n}$ in the bucket.
After bucket filling, the V-turn-2 subtask is performed. It ends with the wheel loader approaching the receiver located at a position $\bm{x}_{n}^{\text{dump}}$, which is the corresponding action parameter $\bm{a}_{n}^{\text{V2}}$. The final subtask is to empty the bucket’s contents onto the load receiver with an action $\bm{a}_{n}^{\text{dump}}$. When the receiver is completely filled, this subtask might require some adjustment action to prevent spillage. A completed loading cycle thus involves a selection of action parameters that we collect in a vector $\bm{a}_{n}=[\bm{a}_{n}^{\text{V1}},\bm{a}_{n}^{\text{load}},\bm{a}_{n}^{\text{V2}},\bm{a}_{n}^{\text{dump}}]$.
A loading cycle has a duration $T_{n}$, requires some amount of work $W_{n}$, and results in a certain amount of mass $M_{n}$ being removed from the pile and added to the receiver. Each of the subtasks contributes to the cycle time and work, while it is only the loading subtask that produces a mass measurement. We attribute each loading cycle a performance, which we measure by the performance vector
@@
It is not a priori clear what is the computational complexity of solving the posed optimization problem. An action sequence $\{\bm{a}_{n}\}_{n=1}^{N}$ renders a sequence of pile states $\{\bm{H}_{n}\}_{n=1}^{N}$ and performance measurements $\{\bm{\mathcal{P}}_{n}\}_{n=1}^{N}$. These cannot be computed independently due to the strong dependency on the evolving pile state. The computational complexity depends ultimately on the dimensionality of the action space $\text{dim}(\bm{a}_{n})$ and on the planning horizon $N$.
@@ 3.1 Assumptions and delimitations
A number of delimitations and simplifying assumptions are made to bring down the difficulty of the problem. We focus on loading a single type of non-cohesive and homogeneous soil, namely gravel. The surrounding ground is assumed to be flat. No soil is spilled from the bucket during the V-turn-2 subtask or when dumped on the receiver. Spillage would affect subsequent performance, either by loss in control precision or traction or by the need to clear the ground.
@@
Automatic bucket filling starts with the loader heading towards a dig location along the edge of the pile at a nominal target speed. To limit the dimensionality of the problem, the heading is always normal to the pile contour, and the edge is discretized into a finite number of candidates of dig locations separated by a step size $\Delta x$.
@@ 4.1 Loading prediction
The dependency on the pile shape is limited to the local pile shape, $\bm{h}$, in the area of the dig location and with a selected heading. Therefore, the function $\bm{\Phi}$ first does a cutout function of the heightmap of the local pile surface. This is then fed as input, along with the loading action parameters, to a deep neural network $\bm{\phi}$. A replace function then substitutes the predicted local heightmap $\bm{h}^{\prime}=\bm{\phi}(\bm{h},\bm{a}^{\text{load}})$ in $\bm{H}$ to obtain the predicted global heightmap $\bm{H}^{\prime}$. Similarly, the loading performance is predicted using a deep neural network, $\bm{\mathcal{P}}^{\text{load}}=\psi(\bm{h},\bm{a}^{\text{load}})$, after first cutting out the local heightmap at the dig location and selected heading. The model has three convolutional layers to encode $\bm{h}$ before being fed with $\bm{a}^{\text{load}}$ to a multi-layer perceptron. The swish activation is used for all layers. The pile state and performance predictor models are illustrated in Fig. 2 and Fig. 3, respectively.
@@ 4.1.1 Optimal loading action parameters
where $\eta$ is the step length and $\nabla=\partial/\partial\bm{a}^{\text{load}}$. As an initial guess, we use $\bm{a}^{\text{load}}_{n}=[1,0,0.0,0.5,0.5]^{\mathsf{T}}$, which corresponds to the machine thrusting deep into the pile, filling the bucket with little or no lifting of it. This type of action has been suggested for uniformly sloped piles of dry homogeneous soil [5], and we expect the optimization process to converge faster with this initialization. The maximum number of iterations is set to 30 with early stopping (patience 3 for tolerance $10^{-4}$). The gradient with respect to $\bm{a}^{\text{load}}$ is calculated by using pytorch.autograd and exploiting that $\psi$ has two input branches, separating $\bm{h}$ and $\bm{a}^{\text{load}}$ so that the encoding part only needs to be evaluated once during the optimization process.
4.2 V-turn prediction
We estimate the transportation time and work consumption between the load receiver $\bm{x}^{\text{dump}}$ and a dig location $\bm{x}^{\text{dig}}$ by numerical integration of an assumed equation of motion for the wheel loader along the two V-turn paths involved in each loading cycle. For generating the V-turn paths, we use cubic B-splines [17]. Each V-turn path is formed by the combination of two spline curves. V-turn-1 is composed of the splines connecting $\bm{x}^{\text{dump}}\to\bm{x}^{\text{V1}}$ and $\bm{x}^{\text{V1}}\to\bm{x}^{\text{dig}}$ with condition that the ingoing and outgoing directions should match in the switch back point $\bm{x}^{\text{V1}}$. Similarly, V-turn-2 is composed of two splines connecting $\bm{x}^{\text{dig}}\to\bm{x}^{\text{V2}}$ and $\bm{x}^{\text{V2}}\to\bm{x}^{\text{dump}}$. See Fig. 4 for an illustration.
@@ 4.2 V-turn prediction
$\left[M_{\text{vehicle}}+M_{\text{load}}\right]\cfrac{dv}{dt}+C_{\text{r}}v=f,$ (12)
@@ 4.4 Look-ahead tree search
where $\bar{\bm{x}}_{k}$ and $\bar{\bm{a}}_{k}$ are greedy choices given the pile state $\bar{\bm{H}}_{k}$, that is, those that minimize $\bm{w}^{\mathsf{T}}\bm{\mathcal{P}}_{k}(\bar{\bm{H}}_{k},\bar{\bm{x}}_{k},\bar{\bm{a}}_{k})$. The pile $\bar{\bm{H}}_{k}$ is, in turn, the result of an expansion from the pile state at the previous search depth, $\bar{\bm{H}}_{k-1}$, using a greedy choice at that level, too [^1] If there is a maximum number loading cycles, $N$, then the summation in Eq. (4.4) is limited by this in case $n+d>N$. Also, if $d=1$, there is no second or higher term on the right-hand side.. Evaluating Eq. (4.4) requires that pile state at the top level pile $\bm{H}_{n}$ is expanded into all the candidate piles $\{\bm{H}_{i}^{\prime}\}_{i=1}^{I}$ using $\{\bm{x}_{i}^{\text{dig}}\}_{i=1}^{I}$ and $\{\bm{a}_{i}^{\text{load}}\}_{i=1}^{I}$. The algorithm is displayed in Alg. 1 and pictorially illustrated in Fig. 1.

### v2

# Optimizing wheel loader performance — an end-to-end approach

@@ Abstract
Wheel loaders in mines and construction sites repeatedly load soil from a pile to load receivers. This task presents a challenging optimization problem since each loading’s performance depends on the pile state, which depends on previous loadings. We investigate an end-to-end optimization approach considering future loading outcomes and transportation costs between the pile and load receivers. To predict the evolution of the pile state and the loading performance, we use world models that leverage deep neural networks trained on numerous simulated loading cycles. A look-ahead tree search optimizes the sequence of loading actions by evaluating the performance of thousands of action candidates, which expand into subsequent action candidates under the predicted pile states recursively. Test results demonstrate that, over a horizon of 15 sequential loadings, the look-ahead tree search is 6 % more efficient than a greedy strategy, which always selects the action that maximizes the current single loading performance, and 14 % more efficient than using a fixed loading controller optimized for the nominal case.
@@
[…]
@@ 1 Introduction
The task of most construction and mining machines, such as wheel loaders, is to perform long sequences of earthmoving operations in dynamic environments. Optimizing this task involves selecting loading and transportation actions that maximize the gain over time, according to some performance measure. This is challenging for several reasons. First, the performance of a single loading cycle depends highly on the local shape of the pile and the soil properties [1]. The loading actions need to be carefully adapted to these conditions. Second, each loading alters the pile shape. From the perspective of future loadings, such a change can be either for the better or worse. There might be actions that are of low performance in the short term but are necessary to achieve higher gains in the future [2]. Therefore, a greedy strategy of always choosing the loading action that maximizes the performance for a single loading might be sub-optimal in the long run. Such a strategy might gradually deteriorate the pile state or lead to excessive transportation between the dig location and the load receiver. This motivates considering future states when selecting the actions.
@@
This paper investigates methods for solving the wheel loader end-to-end optimization problem and analyzes the obtained solutions. The research questions include what characterizes the optimal action sequence, how sensitive the results are to the initial pile state and planning horizon, what the computational demands are, and the feasibility of practical use of the method. The idea is that the methods could be implemented in site management systems with high-precision monitoring of equipment and the 3D topography, sending coordinated work plans to the individual machines.
We use a look-ahead tree search approach illustrated in Fig. 1. Future action candidates for the foreseen future states are identified recursively. The action candidates are evaluated, using data-driven models developed in previous research [3], to compute their respective performance and expand into further action candidates under the new states. This approach is, reportedly, similar to how experienced operators intuitively plan their work, but we expect that a computerized model and search algorithms can make more accurate predictions and plan over longer horizons. To understand the importance of the look-ahead search, we examined the effect of using different planning horizons and objectives.
@@
The wheel loader is assumed to have a high-level planner and a low-level control system, which work as the hierarchical architecture described in refs. [4] and [5]. The high-level planner selects a subtask action, such as bucket filling, dumping, or V-shaped transportation, according to the current state of the machine and the work site. A low-level controller performs the subtasks given a motion plan and control parameters. For bucket filling, the machine is equipped with a type of admittance controller [6]. Our controller takes four action parameters that determine how reactive the boom and bucket actuation is in response to the momentaneous dig force [3]. The performance of a loading cycle is measured in terms of mass, time, and work, with contributions from each subtask.
@@ 2 Related work
What is a good dig location in the pile, given its shape in terms of a heightmap, has been studied from the perspective of both bucket filling and the transportation between the dig and the dump location. In [7], a coarse-level planner finds the points along the pile contour closest to the load receiver, and a refined planner was used to avoid curved regions of the pile surface to minimize bucket-side loading and drop-in filling efficiency. A sequential coarse-to-fine planner was studied in [8] and compared to a greedy planner. The coarse planner first produces a sequence of dig regions ordered radially around the pile centroid. The fine planner then selects the dig points inside each region that have the best local pile shape. Simulations (cellular automata) of 50 consecutive loadings showed that the coarse-to-fine strategy maintains a good pile shape and performs at 80–90 % while a greedy strategy (always selecting the dig location with optimal pile shape) initially performs equal or better for ten loadings but drops down to about 60 % after 25 loadings.
@@
The optimal bucket filling trajectory was researched in [18] using the discrete element method (DEM) and used in [19] to find the optimal short loading cycle using dynamic programming. For the sake of computational efficiency, the simulations were limited to quasi-2D piles sloped with the material’s angle of repose. The effect of the pile shape on the optimal trajectory or the resulting 3D pile state has not been explored using DEM simulation, apart from the work in [10].
@@ 3 Problem formulation
We focus on the short loading cycle, in which the wheel loader repeatedly loads and dumps soil from a pile to a load receiver. Each cycle can be divided into a sequence of subtasks: V-turn-1, loading, V-turn-2, and dumping. These subtasks are illustrated in Fig. 1. We assume the wheel loader is equipped with a low-level control system for each subtask. A high-level system (operator or agent) activates the low-level systems by selecting a set of subtask action parameters: $\bm{a}^{\text{V1}}$, $\bm{a}^{\text{load}}$, $\bm{a}^{\text{V2}}$, and $\bm{a}^{\text{dump}}$, respectively.
At the beginning of each cycle $n=1,2,\ldots,N$, the wheel loader is located by the load receiver at a position $\bm{x}_{n}^{\text{dump}}$. The V-turn-1 subtask is to drive the wheel loader to a selected loading position $\bm{x}_{n}^{\text{dig}}$ somewhere along the edge of the nearby pile. A low-level controller plans and executes a V-turn motion that starts at $\bm{x}_{n}^{\text{dump}}$ and ends at $\bm{x}_{n}^{\text{dig}}$, avoiding collisions with the environment, and with endpoints heading normal to the load receiver and the pile, respectively. As subtask action parameters, we use the target loading position, $\bm{a}_{n}^{\text{V1}}=\bm{x}_{n}^{\text{dig}}$.
The next subtask is the actual loading from the pile. It starts with the machine approaching the selected loading position and ends with the vehicle reversing from the pile with a filled bucket. The task is carried out by an automatic bucket filling controller with some action parameters, $\bm{a}_{n}^{\text{load}}$, that may be adapted to the current pile shape and material properties. The loading transforms the pile from a state $\bm{H}_{n}$ into a new state $\bm{H}_{n+1}$ and results in some amount of soil mass $M_{n}$ in the bucket. We identify the state with the geometric shape of the pile.
After bucket filling, the V-turn-2 subtask is performed. It ends with the wheel loader approaching the receiver located at a position $\bm{x}_{n}^{\text{dump}}$, which is the corresponding action parameter $\bm{a}_{n}^{\text{V2}}$. The final subtask is to empty the bucket’s contents onto the load receiver with an action $\bm{a}_{n}^{\text{dump}}$. When the receiver is filled, this subtask might require some adjustment action to prevent spillage. A completed loading cycle thus involves a selection of action parameters that we collect in a vector $\bm{a}_{n}=[\bm{a}_{n}^{\text{V1}},\bm{a}_{n}^{\text{load}},\bm{a}_{n}^{\text{V2}},\bm{a}_{n}^{\text{dump}}]$.
A loading cycle involves spending some amount of mechanical work $W_{n}$ and time $T_{n}$ to displace a certain mass $M_{n}$ from the pile to the receiver. Each of the subtasks contributes to the cycle time and work, while it is only the loading subtask that produces a mass measurement. We attribute each loading cycle a performance, which we measure by the performance vector
@@
The posed optimization problem is computationally challenging to solve. An action sequence $\{\bm{a}_{n}\}_{n=1}^{N}$ renders a sequence of pile states $\{\bm{H}_{n}\}_{n=1}^{N}$ and performance measurements $\{\bm{\mathcal{P}}_{n}\}_{n=1}^{N}$. These cannot be computed independently due to the strong dependency on the evolving pile state. The computational complexity for exhaustive search scales exponentially as $(t_{\bm{a}}D_{\bm{a}})^{N}$, where $D_{\bm{a}}$ is the number of action candidates that must be evaluated for each pile state and $t_{\bm{a}}$ is the computational time for doing so.
@@ 3.1 Assumptions and delimitations
A number of delimitations and simplifying assumptions are made to bring down the difficulty of the problem. We focus on loading a single type of non-cohesive and homogeneous soil, namely gravel. The surrounding ground is assumed to be flat. We assume no soil is spilled from the bucket during the V-turn-2 subtask or when dumped on the receiver. Spillage would affect subsequent performance, either by loss in control precision or traction or by the need to clear the ground.
@@
Automatic bucket filling starts with the loader heading towards a dig location along the edge of the pile at a nominal target speed. To limit the dimensionality of the problem, the heading is always normal to the pile contour, and the edge is discretized into a finite number of candidates of dig locations separated by a step size smaller than the width of a bucket.
@@ 4.1 Loading prediction
The dependency on the pile shape is limited to the local pile shape, $\bm{h}$, in the area of the dig location and with the selected heading. Therefore, the function $\bm{\Phi}$ first does a cutout operation of the global heightmap to obtain the local pile surface. This is then fed as input, along with the loading action parameters, to a deep neural network $\bm{\phi}$. A replace function then substitutes the predicted local heightmap $\bm{h}^{\prime}=\bm{\phi}(\bm{h},\bm{a}^{\text{load}})$ into $\bm{H}$ to obtain the predicted global heightmap $\bm{H}^{\prime}$. Similarly, the loading performance is predicted using a deep neural network, $\bm{\mathcal{P}}^{\text{load}}=\psi(\bm{h},\bm{a}^{\text{load}})$, after first cutting out the local heightmap at the dig location and selected heading. The model has three convolutional layers to encode $\bm{h}$ before being fed with $\bm{a}^{\text{load}}$ to a multi-layer perceptron. The swish activation is used for all layers. The pile state and performance predictor models are illustrated in Fig. 2 and Fig. 3, respectively.
@@ 4.1.1 Optimal loading action parameters
where $\eta$ is the step length and $\nabla=\partial/\partial\bm{a}^{\text{load}}$. As an initial guess, we use $\bm{a}^{\text{load}}_{n}=[1,0,0.0,0.5,0.5]^{\mathsf{T}}$, which corresponds to the machine thrusting deep into the pile, filling the bucket with little or no lifting of it [3]. This type of action has been suggested for uniformly sloped piles of dry homogeneous soil [20], and we expect the optimization process to converge faster with this initialization. The maximum number of iterations is set to 30 with early stopping (patience 3 for tolerance $10^{-4}$). The gradient with respect to $\bm{a}^{\text{load}}$ is calculated by using pytorch.autograd and exploiting that $\psi$ has two input branches, separating $\bm{h}$ and $\bm{a}^{\text{load}}$ so that the encoding part only needs to be evaluated once during the optimization process.
4.2 V-turn model
A loading cycle involves moving the wheel loader back and forth between the pile and the load receiver in a V-shaped pattern connecting the dump location $\bm{x}^{\text{dump}}$ and dig location $\bm{x}^{\text{dig}}$. We estimate the transportation time and work by numerical integration of an assumed equation of motion for the wheel loader, including its motor strength, variable mass, and rolling resistance. For generating the V-turn paths, we use cubic B-splines [21]. Each V-turn path is formed by the combination of two spline curves. V-turn-1 is composed of the splines connecting $\bm{x}^{\text{dump}}\to\bm{x}^{\text{V1}}$ and $\bm{x}^{\text{V1}}\to\bm{x}^{\text{dig}}$ with condition that the ingoing and outgoing directions should match in the switch back point $\bm{x}^{\text{V1}}$. Similarly, V-turn-2 is composed of two splines connecting $\bm{x}^{\text{dig}}\to\bm{x}^{\text{V2}}$ and $\bm{x}^{\text{V2}}\to\bm{x}^{\text{dump}}$. See Fig. 4 for an illustration.
@@ 4.2 V-turn model
$M_{\text{tot}}\cfrac{dv}{dt}+C_{\text{r}}v=f,$ (12)
@@ 4.4 Look-ahead tree search
where $\bar{\bm{x}}_{k}$ and $\bar{\bm{a}}_{k}$ are greedy choices given the pile state $\bar{\bm{H}}_{k}$, that is, those that minimize $\bm{w}^{\mathsf{T}}\bm{\mathcal{P}}_{k}(\bar{\bm{H}}_{k},\bar{\bm{x}}_{k},\bar{\bm{a}}_{k})$. The pile $\bar{\bm{H}}_{k}$ is, in turn, the result of an expansion from the pile state at the previous search depth, $\bar{\bm{H}}_{k-1}$, using a greedy choice at that level too [^1] If there is a maximum number loading cycles, $N$, then the summation in Eq. (4.4) is limited by this in case $n+d>N$. Also, if $d=1$, there is no second or higher term on the right-hand side.. Evaluating Eq. (4.4) requires that pile state at the top level pile $\bm{H}_{n}$ is expanded into all the candidate piles $\{\bm{H}_{i}^{\prime}\}_{i=1}^{I}$ using $\{\bm{x}_{i}^{\text{dig}}\}_{i=1}^{I}$ and $\{\bm{a}_{i}^{\text{load}}\}_{i=1}^{I}$. The algorithm is displayed in Alg. 1 and pictorially illustrated in Fig. 1.